OpenStudy (anonymous):

Calc 3 question, please help! Iterated Integrals

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi}\int\limits_{1}^{2}ysin(xy)dxdy\]

OpenStudy (anonymous):

I need help integrating with respect to x

OpenStudy (anonymous):

\[\begin{align*}\int\limits_{0}^{\pi}\int\limits_{1}^{2}y\sin(xy)~dx~dy&=\int\limits_{0}^{\pi}\int\limits_{y}^{2y}y\sin u~\left(\frac{du}{y}\right)~dy&\text{set }u=xy\\\\ &=\int\limits_{0}^{\pi}\int\limits_{y}^{2y}\sin u~du~dy\\\\ &=-\int\limits_{0}^{\pi}[\cos u]_y^{2y}~dy\\\\ &=-\int\limits_{0}^{\pi}(\cos2y-\cos y)~dy\\\\ &=\int\limits_{0}^{\pi}(\cos y-\cos2y)~dy \end{align*}\]